OpenStudy (czarluc):

y=(10^x-10^-x)/(10^x+10^-x) find x in terms of y

hartnn (hartnn):

try Componendo-Dividendo, If you know what that is

hartnn (hartnn):

If a/b = c/d by componendo-Dividendo, (a+b)/(a-b) = (c+d)/(c-d)